OpenStudy (anonymous):

cos theta=1/9 what is sin of theta?

OpenStudy (ajprincess):

R u familiar with this formula \(\cos^2\theta+\sin^2\theta=1\) @Holly.hamilton1?

OpenStudy (anonymous):



OpenStudy (anonymous):

so how would i find sin?